Prove that 2 1 n 2 n 2
n n 1
> + ∀ >

, using the inequalities of Unit 6.
1
Expert's answer
2019-10-01T09:54:23-0400

Well, in general case:


"2n^2>n^2-n-1.""2n^2-n^2+n+1>0.""n^2+n+1>0."


Now we have a quadratic inequality. The quadratic equation


"n^2+n+1=0\\"

has the following discriminant:


"D=1^2-4*1*1=-3\\"

"D<0", therefore quadratic equation has no real roots (only complex ones). As far as the coefficient before the second power is positive ("a=1"), parabola, which represents equation, is oriented upward, thereby confirming inequality "\, n^2+n+1>0" for "\forall" "n\in\real".


In particular cases of integers:

  1. For "n\geq0" inequality "n^2+n+1>0" is obviously true.
  2. For "n<0". Let "k=-n>0". Then we can rewrite inequality:

"k^2-k+1>0\\"

"k^2-k=k(k-1)>0" for any "k>1.\\"

Thus, inequality "n^2+n+1>0" is true for "\forall" "n\in\mathbb{Z}" .